OpenStudy (anonymous):

L'Hopital's lim x->-inf (1-(3/x))^(2x) Im stuck at lim """ 2xln((x-3)/x))

OpenStudy (amistre64):

change 2x to (2x)^(-1) maybe to make this a fraction

OpenStudy (roadjester):

Is this \[\lim_{x \rightarrow -\infty}{\frac 1{3x}}^{2x}\]?

OpenStudy (amistre64):

\[\lim \frac{ln(1-\frac{3}{x})}{1/2x}\]

OpenStudy (anonymous):

how do you explain that mathematically? divide by 2x^2?

OpenStudy (amistre64):

\[\frac{1}{a}=a^{-1}\]

OpenStudy (anonymous):

that gives you 0/(1/0) does lhop work for that?

OpenStudy (anonymous):

1/inf*

OpenStudy (amistre64):

its indeterminant :) but id have to chk to be sure

OpenStudy (roadjester):

0/(1/0) is 0/infinity

OpenStudy (roadjester):

L'hopital's only works for 0/0 or infinity/infinity

OpenStudy (anonymous):

yeah so how do I know to stop there? I mean couldn't we manipulate it to get 0/0 or inf/inf?

OpenStudy (roadjester):

Yes, what Amistre64 wrote does get you that

OpenStudy (roadjester):

3/infinity is 0 ln(1)=0

OpenStudy (amistre64):

1-3/inf = 1; ln(1) = 0

OpenStudy (anonymous):

I get that part but can't I do something else to make it so its 0/0 or something

OpenStudy (roadjester):

It is already 0/0

OpenStudy (anonymous):

with what amistre wrote we have 0/ (1/-inf)) right?

OpenStudy (amistre64):

i spose you can do anything you want to it; but why would some other thing have to be worked out if we got something that works now?

OpenStudy (roadjester):

yes

OpenStudy (amistre64):

1/2x = 1/inf = 0

OpenStudy (anonymous):

I was under the impression that we determined that we can't use l'hops rule.

OpenStudy (roadjester):

we can

OpenStudy (roadjester):

Let me write it out shall I?

OpenStudy (anonymous):

so it is 0/0 ahh, my apologies calc teachers here don't teach us anything very useful like we can use 1/inf lol

OpenStudy (amistre64):

lim ln(1-3/-inf) -0 x>-inf --------- = ---- if anything :) 1/2(-inf) -0

OpenStudy (anonymous):

this is a huge mess..

OpenStudy (roadjester):